### v2

# Sharing Matters: Analysing Neurons Across Languages and Tasks in LLMs

Abstract
Large language models (LLMs) have revolutionized the field of natural language processing (NLP), and recent studies have aimed to understand their underlying mechanisms. However, most of this research is conducted within a monolingual setting, primarily focusing on English. Few studies attempt to explore the internal workings of LLMs in multilingual settings. In this study, we aim to fill the research gap by examining how neuron activation is shared across tasks and languages. We classify neurons into four distinct categories based on their responses to a specific input across different languages: all-shared, partial-shared, specific, and non-activated. This categorization is combined with a study of neuron attribution, i.e. the importance of a neuron w.r.t an output. Our analysis reveals the following insights: (i) the patterns of neuron sharing are significantly affected by the characteristics of tasks and examples; (ii) neuron sharing does not fully correspond with language similarity; (iii) shared neurons play a vital role in generating responses, especially those shared across all languages. These findings shed light on the internal workings of multilingual LLMs and pave the way to the future research. We will release the code to foster research in this area.
Sharing Matters: Analysing Neurons Across Languages and Tasks in LLMs
Weixuan Wang1, Barry Haddow1, Minghao Wu2, Wei Peng3, Alexandra Birch1 1School of Informatics, University of Edinburgh 2Monash University 3Huawei Technologies Co., Ltd. {weixuan.wang, bhaddow, a.birch}@ed.ac.uk minghao.wu@monash.edu peng.wei1@huawei.com
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in recent studies, excelling in both understanding and generating text across various languages (OpenAI, 2023; Zhang et al., 2023; Zhao et al., 2024). Despite their proven effectiveness, the intricate mechanisms underlying their processing remain largely opaque. This opacity has given rise to a growing field of research aimed at interpreting the internal workings of the Transformer architecture (Elhage et al., 2021; Yu et al., 2023). To enhance interpretability and investigate specific aspects of model behavior, researchers have increasingly focused on the components of these models. Recent studies have explored the role of Feed-Forward Networks (FFNs) within LLMs, proposing that these components function as key-value memories for storing factual and linguistic knowledge (Geva et al., 2020, 2022; Ferrando et al., 2023). While these studies have analyzed neuron behaviors based on activation states in monolingual settings, there remains a significant gap in our understanding of how neurons behave in multilingual contexts.
Significant discrepancies have been observed in neuron activation patterns across languages, despite identical semantic outputs, as highlighted in Figure 1. This phenomenon raises critical questions about how LLMs manage multilingual representations at the neuronal level. In a similar line of reasoning, Bhattacharya and Bojar (2023) categorized neurons into two coarse-grained groups: language-agnostic (shared across languages) and language-specific (unique to a language). However this categorization oversimplifies the complexity observed in cross-lingual studies, where neuron overlap varies significantly between languages (Stanczak et al., 2022; Zhao et al., 2023; Liu et al., 2024). Additionally, most research has been confined to single-task analyses, overlooking how neuron types might shift across diverse tasks (Bhattacharya and Bojar, 2023; Tang et al., 2024; Tan et al., 2024). This underscores the need for a more nuanced, fine-grained classification method to enhance our understanding of the multifaceted roles of neurons in multilingual LLMs.
To bridge this gap, our research introduces a fine-grained classification of neurons, enabling a detailed exploration of their roles across languages and tasks. We systematically analyse neuron behaviors across three distinct tasks in ten languages, including reasoning, fact probing, and question answering tasks, using various diverse model backbones: BLOOMZ-7b, LLaMA2-7b-chat, BLOOM-7b, XGLM, and mT0. For a specific input and its corresponding translations in 10 languages, we reformulate the activation state of neurons to four distinctive types to represent multilingual neuron roles (see Section 3). All-shared neurons are neurons that remain active for all the translations regardless of language. Partial-shared neurons are activated only for translations in certain languages. Specific neurons are activated exclusively for translations in one language. Non-activated neurons are not activated at all for any the input’s translations. We then use the ratio of neurons in each type to analyze the activation pattern across various tasks. Furthermore, we evaluate the importance of each type of neurons in predicting answers using contribution scores and effective scores (Geva et al., 2022; Voita et al., 2023). Finally, we confirm our analysis through experiments on downstream tasks.
We provide substantial empirical evidence detailing neuron activation patterns and attributions in this study, leading to several significant findings. Here are the main takeaways:
Neuron activation patterns are considerably task-specific and example-specific. We observe the ratio of four types of neurons is related to the task (see Section 5.1), where the neuron activation pattern of XNLI is significantly different from that of Fact Probing. We also demonstrate that the neuron types commonly vary with different test examples (see Section 5.2).
Similar languages do not always lead to high degree of neuron sharing. Our empirical results show that the languages from the same language family, such as German and French, do not exhibit higher degree of neuron sharing, compared to the languages from distinct language families, such as German and Chinese (see Section 5.3).
Shared neurons are crucial in generating responses. Both the contribution scores and the effective scores highlight the significance of all-shared neurons (see Section 6). Our experiments further reveal that disabling the all-shared neurons can lead to a substantial decrease (up to 91.6%) in model performance on downstream tasks (see Section 7).
2 Related Work
The black-box nature of LLMs has given rise to an area of research which aims to interpret the internal mechanism of the Transformer architecture (Elhage et al., 2021; Yu et al., 2023). More recently, several studies on LLMs have advanced our understanding of how neurons acquire task-specific knowledge. For instance, Ferrando et al. (2023); Dai et al. (2022); Geva et al. (2020, 2022) investigated how FFN blocks function as key-value memories and proved that factual knowledge is stored in the neurons. Research work on the sparsity of neurons in FFN blocks showed that many neurons are inactive in various tasks (Zhang et al., 2022; Li et al., 2023). Voita et al. (2023) located these “dead” neurons in the lower part of the model (close to inputs) in the English scenario. Despite the insights obtained, these studies have focused exclusively on a monolingual setting.
For multilingual neuron analysis, Bhattacharya and Bojar (2023) explored the neuron sharing between two languages. Tang et al. (2024); Tan et al. (2024); Liu et al. (2024); Kojima et al. (2024) classified neurons in an FFN block to language-specific and language-agnostic based on predefined threshold, assuming that a neuron’s classified type remained consistent across different examples. However, the broad classification into two groups is inadequate for detailed multilingual analysis. Additionally, these studies classified neurons based on the single task (Tan et al., 2024; Liu et al., 2024), without considering the potential adaptation of neurons under various languages and semantics brought forth by inputs from various multilingual tasks. We investigate neurons’ behaviors across multiple languages and tasks to this end.
3 Neuron Classification
In this section, we provide a detailed description of the 4-way neuron classification that we propose. We begin with some background concerning neurons in the FFN block (Section 3.1). Following this, we define the four types of neurons (Section 3.2), and explain how we collect activation patterns by averaging the ratio of each neuron type across the entire test set (Section 3.3).
3.1 Neurons in FFN Blocks
A neuron inside the FFNs is defined as a linear transformation of an input representation followed by a non-linear activation (Tang et al., 2024). Every feed-forward block at layer involves two linear transformations separated by a point-wise activation function. Biases are omitted for the sake of clarity:
where , are linear parameter matrices, and is a non-linear activation function, where rows in and columns in are viewed as -dimensional keys and values , respectively. We define the behavior of a neuron to be the output of the function immediately after the elementwise nonlinearity. is the count of neurons. And the output of neurons determines the weighting of the corresponding values in .
For the -th neuron and corresponding key , value and activation value , we can express this relationship using the following formulation:
When such a neuron is activated , so it updates the residual stream by pulling out the corresponding value . If a neuron is inactive, the activation value is .
3.2 Definition of Four Types of Neurons
In this work, we categorize the neurons into four types based on their activation values and detail the neuron classification in this section. To ablate the impact of semantic discrepancies across languages, the datasets used in this work are initially in English and then translated into foreign languages (see Section 4.1), so we can formulate the -th example as , where indicates the -th language and is the total number of languages.
Given the -th example , the set of all-shared neurons at the -th layer can be defined as:
where is the set of all the neurons at the -th layer and is the -th neuron in . Furthermore, the non-activated neurons is the set of neurons whose activation value less than or equal to zero in all languages, as follows:
Moreover, the specific neurons are the neurons only activated in one specific language and not activated in any other languages, defined as follows:
Lastly, the remaining neurons are partial-shared neurons as they are activated by inputs from a subset of languages:
Note that, unlike Bhattacharya and Bojar (2023); Tang et al. (2024) who consider the sub-word activation states and thus capture incomplete semantics, we only examine the activation state of the last token, as that is when the LLM performs the prediction task.
3.3 Collecting Neuron Activation Patterns
In this section, we explain how to quantify neuron activation patterns. Specifically, we measure the percentage of each type of neuron relative to the total number of neurons. Given the -th test instance, the percentage of each neuron type at the -th layer can be defined as follows:
where denotes the number of elements in the set. Consequently, the aggregated neuron activation pattern for one dataset contain instances at the -th layer can be defined as:
4 Experimental Setting
4.1 Multilingual Tasks
We perform analysis on neurons in FFN blocks of various LLMs, harnessing their multilingual capabilities in three diverse tasks which consist of multilingual parallel sentences, including XNLI (Conneau et al., 2018), Fact Probing (Fierro and Søgaard, 2022), and Cross-lingual Knowledge Editing (KE) (Wang et al., 2023). For the Cross-lingual KE, we analyze the LLMs in two setups, including EN (Edit) ALL (Test) and ALL (Edit) EN (Test). More details are described in Appendix A.2.
These tasks cover 10 diverse languages, including English (en), German (de), Spanish (es), French (fr), Portuguese (pt), Russian (ru), Thai (th), Turkish (tr), Vietnamese (vi), and Chinese (zh). Prompts are detailed in Appendix A.3.
4.2 Base LLMs
We mainly analyze the behaviors and attributions of neurons in an instruction-finetuned model BLOOMZ-7b (Muennighoff et al., 2023). We also include the analysis of other decoder-only models: BLOOM-7b (Scao et al., 2022), LLAMA2-7b-chat (Touvron et al., 2023), XGLM (Lin et al., 2022), and an encoder-decoder model mT0 (Muennighoff et al., 2023) in the Appendix A.4.1.
5 Behaviors of Four Types of Neurons
In this section, we present our in-depth analysis with regard to the neuron activation across task (Section 5.1), examples within the same task (Section 5.2), and language (Section 5.3).
For each task we use parallel test texts in ten languages as input, and record the type of each neuron. Subsequently, we calculate and compare the percentage of the four types of neurons compared to the total neurons. We use BLOOMZ-7b as the backbone LLM to investigate the behaviors of four types of neurons, and the results of three tasks are shown in Figure 3 respectively.
5.1 Neuron Activation Across Tasks
The neuron activation pattern is strongly related to the task performed.
As shown in Figure 3, non-activated neurons are more prevalent overall than other types. However, the pattern observed in the Fact Probing task differs significantly from the other two tasks. In the Fact Probing task, there are more partial-shared neurons and specific neurons, whereas the other tasks involve far more all-shared neurons. In LABEL:bloom-llama-3tasks, the activation patterns of the BLOOM-7b and LLaMA2-7b-chat LLMs are presented, highlighting variability across models while affirming the consistent influence of the task on neuron activation patterns. Additionally, a comparison between BLOOM and BLOOMZ LLMs indicates that instruction fine-tuning in BLOOMZ appears to foster a higher proportion of all-shared neurons, suggesting better alignment across multilingual representations.
5.2 Neuron Activation Across Examples
The type of a neuron varies across examples within the same task.
As shown in Figure 3, we have analysed the percentage of each type of neurons across tasks. In this experiment, we now investigate how neurons behave across all examples within one task, aiming to determine whether neurons categorized as a specific type retain their classification across different semantics. Take the Fact Probing task as an example, the results are shown in Figure 5 and in Appendix A.4.2 (for other tasks). To our surprise, almost no neuron (identified by its index) retains its type (in our four-ways classification) across all examples. This finding contradicts prior research that suggested neuronal classification remains unchanged amidst variation in input semantics. Our study thus underscores that the classified type of a neuron is sensitive to the semantics of the encountered inputs, even within the same task.
Non-activated neurons do not stay inactive across examples.
In Figure 5 we observe that less than 10% of non-activated neurons remain inactive across all examples. This small proportion of persistently inactive neurons reflects the distributed nature of knowledge representation in LLMs. That is why we should execute caution when pruning neurons, as we may damage the overall performance of an LLM.
Non-activated neurons do not stay inactive across all tokens in an input.
Expanding our focus, we examine whether non-activated neurons remain inactive across all tokens of an input, differing from our prior analysis which concentrated only on the last token. Specifically, for each input sequence with tokens , we record non-activated neurons of each token, w.r.t the intersection of index. A neuron is counted when it stays inactive for every token of one test input sequence. As shown in Figure 6, fewer than 0.8% non-activated neurons consistently stay inactive for each token of the input, which further highlights the variability in neuron activity.
5.3 Neuron Activation Across Languages
Neuron sharing does not completely align with language similarity.
In exploring the relationship between language similarity and neuron sharing, we specifically analyse the proportion of shared neurons between German and other languages within the category of partial-shared neurons. Additionally, we examine the ratio of specific neurons for each language. These analyses are conducted within the Fact Probing task. As shown in Figure 7, similar languages do not share more neurons; for instance, German and Chinese (in different language families) have a similar ratio of shared neurons to that of German and French (in the same language family) in the upper layers. Furthermore, we observe that there is no consistent pattern with regard to the ratio of specific neurons across the languages studied. We leave the further exploration on this matter to the future work. More results on XNLI task can be found in Appendix A.4.3.
More languages encourage greater neuron sharing.
We also investigate the impact of the number of languages on the percentage of four types of neurons. As illustrated in Figure 8, our study include configurations of 3, 5, 7, and 9 languages on the XNLI task. Results show a noticeable escalation in the percentage of partial-shared neurons as more languages are introduced, indicating that language configuration affects the results of model’s neuron sharing. This analysis emphasizes the potential influence of language selection on multilingual interpretability analysis.
6 Neuron Attributions
In the preceding experiments, we examined the proportions of the four types of neurons across tasks and languages. In this section, we evaluate importance of neurons in answer generation using contribution scores (Section 6.2), and assess the influence of neurons in predicting correct answers through effective scores (Section 6.3) with BLOOMZ-7b backbone LLM. We start with the fundamentals of the contribution analysis method (Section 6.1).
6.1 Contribution Score and Effective Score
In this section, we introduce two measures of quantifying the importance of neurons: Contribution Score and Effective Score.
Contribution Score
Inspired by Geva et al. (2022), in order to judge the importance of neurons in generating answers, we analyze their contributions to the output. The contribution score of the -th neuron at -th layer to an FFN output is:
which is the proportion of its weight to the sum of weights of all neurons in the FFN block. is the absolute value of activation value and is the L2-norm of value .
Effective Score
Following Geva et al. (2022) and Voita et al. (2023), the effective score of the -th neuron at -th layer can be defined as:
where is the embedding of the correct answer . A larger has a higher probability to produce the correct answer , while a negative reduces the probability in generating . In this way, we can quantify the effect of a neuron on the reference answer. We give detailed descriptions about the neuron projection in Appendix A.1.
Comparison
The contribution score quantifies how much neurons contribute to the overall generation process, regardless of the correctness of the output, while the effective score incorporates the correct answer’s embedding to measure the neurons’ impact specifically in generating accurate responses. Although both metrics assess neuronal influence, the key difference lies in their consideration of answer correctness: the contribution score focuses solely on the generation process, whereas the effective score takes into account the neurons’ role in the accuracy of the resulting answer.
6.2 The Contribution of Neuron Types
In this section, we explore the influence of each neuron type using the contribution score described in Section 6.1. We examine the proportion of the four types of neurons among the top 5% contribution score under inputs in each language. The proportions of neurons in the Cross-lingual KE task are depicted in Figure 9 (and the overall results are shown in Figure 21 in Appendix A.5.1).
All-shared neurons are the top contributing neurons to outputs.
Figure 9 depicts a stacked column chart where each segment color represents a specific neuron type’s contribution, collectively totaling 100%. It can be observed that all-shared neurons constitute the largest segment, highlighting their significant role irrespective of language input. Partial-shared neurons are the second most influential, affecting mainly the latter half of the model. It is not surprising that the specific neurons group has limited influence on Cross-lingual KE (EN (Edit) ALL (Test)) outputs as they feature in a particular language type of inputs.
We further analyze the average and total contribution scores of all neurons across four categories. As shown in Figure LABEL:bloomz-contribution_absolute-allspecific_twolangs, all-shared neurons score the highest averages, especially in upper layers, but they constitute only a minor portion of the neuron population (in Figure 3). Meanwhile, non-activated and partial-shared neurons contribute significantly to the total scores in the XNLI and fact-probing tasks respectively, predominantly due to their larger numbers. The disparity in total contributions across tasks confirms our earlier findings in Section 5.1. We leave further exploration regarding to task-specific neuron activation to the future work.
| language | neuron type | max | min | avg | var |
| en | all-shared | 1.85 | -0.94 | 0.07 | 0.36 |
| partial-shared | 0.22 | -0.16 | 1.2e-4 | 1.9e-4 | |
| specific | 0.02 | -0.02 | 2.5e-4 | 3.5e-5 | |
| non-activated | 0.04 | -0.03 | 2.1e-4 | 5.8e-6 | |
| de | all-shared | 1.03 | -0.60 | 0.02 | 0.07 |
| partial-shared | 0.13 | -0.13 | 6.7e-5 | 7.1e-5 | |
| specific | 0.07 | -0.03 | 7.5e-5 | 2.3e-5 | |
| non-activated | 0.02 | -0.01 | 7.0e-5 | 2.9e-6 |
6.3 The Effectiveness of Neuron Types
In this section, we explore the effectiveness of each neuron type using the effective score described in Section 6.1.
All-shared neurons have a greater influence on generating correct answers.
In the Cross-lingual KE (EN (Edit) ALL (Test)) task, we analyze the effectiveness scores (maximum, minimum, average, and variance) for each neuron type throughout all layers of the BLOOMZ model, as shown in Table 1. All-shared neurons display both the highest maximum and the lowest minimum scores, suggesting they are pushed (or eliminated) strongly by the activation functions. By comparison, while all-shared neurons and partial-shared neurons show a wide variance in scores (1.85 vs. -0.94 and 0.22 vs. -0.16, respectively, in English), specific neurons and non-activated neurons have much tighter score ranges ( 0.07). Additionally, we demonstrate the maximum effective scores across layers in Figure 11. It could be observed the neurons in the upper layers play a more crucial role in determining model performance.
| settings | pct. | en | de | es | fr | ru | th | tr | vi | zh |
| XNLI task | ||||||||||
| baseline | 0% | 53.8 | 41.8 | 50.3 | 49.0 | 47.6 | 40.9 | 34.9 | 50.5 | 51.1 |
| w/o. all-shared | 9.85% | 16.7 | 3.5 | 10.1 | 10.0 | 6.6 | 9.0 | 1.4 | 12.1 | 14.5 |
| w/o. partial-shared | 10.30% | 52.9 | 40.4 | 49.7 | 47.6 | 49.2 | 40.3 | 36.1 | 50.0 | 50.0 |
| w/o. specific | 3.12% | 53.7 | 41.7 | 50.3 | 48.9 | 47.4 | 40.6 | 35.3 | 50.4 | 49.3 |
| w/o. non-activated | 76.73% | 36.6 | 31.6 | 33.6 | 33.4 | 29.5 | 31.3 | 28.3 | 34.5 | 23.5 |
| w/o. random | 5% | 53.2 | 42.2 | 50.7 | 48.8 | 47.4 | 40.2 | 34.5 | 50.1 | 50.9 |
| w/o. random | 15% | 53.1 | 41.8 | 50.1 | 48.9 | 47.3 | 40.8 | 33.8 | 50.1 | 50.4 |
| w/o. random | 25% | 52.6 | 41.7 | 50.3 | 48.8 | 46.0 | 38.8 | 36.2 | 50.7 | 49.7 |
| w/o. random | 75% | 36.0 | 28.7 | 40.7 | 36.7 | 28.9 | 25.4 | 23.0 | 38.5 | 32.9 |
| Cross-lingual KE (EN (Edit) ALL (Test)) task | ||||||||||
| baseline | 0% | 96.2 | 55.1 | 48.9 | 49.5 | 30.6 | 6.3 | 38.8 | 51.7 | 33.4 |
| w/o. all-shared | 8.69% | 11.0 | 4.9 | 6.1 | 4.9 | 1.9 | 0.4 | 1.5 | 6.1 | 2.9 |
| w/o. partial-shared | 13.20% | 90.2 | 51.7 | 48.9 | 48.9 | 25.4 | 5.5 | 35.5 | 50.9 | 38.4 |
| w/o. specific | 4.87% | 96.1 | 54.4 | 48.7 | 48.9 | 30.4 | 6.3 | 37.9 | 51.7 | 28.5 |
| w/o. non-activated | 73.20% | 36.1 | 15.8 | 18.2 | 17.8 | 1.9 | 9.8 | 19.9 | 10.6 | 16.3 |
| w/o. random | 5% | 96.1 | 46.9 | 36.6 | 40.4 | 0.8 | 4.4 | 28.7 | 40.8 | 10.1 |
| w/o. random | 15% | 94.8 | 47.1 | 36.1 | 39.4 | 0.8 | 4.3 | 28.4 | 40.4 | 11.1 |
| w/o. random | 25% | 91.5 | 46.8 | 36.2 | 38.6 | 1.1 | 4.4 | 27.9 | 40.4 | 12.1 |
| w/o. random | 75% | 11.1 | 5.5 | 9.3 | 11.3 | 0.1 | 2.7 | 1.9 | 8.9 | 7.1 |
7 Verification on Downstream Tasks
We analyse the behaviors of different neuron types in Section 5 and explore the contribution and effectiveness of various types of neurons in Section 6, demonstrating that neuron shared across languages and tasks are of vital importance from an internal representation perspective. In this section, we conduct experiments on downstream tasks to verify our analytical findings regarding the significance of neuron sharing.
We assess accuracy changes by selectively deactivating four specific types of neurons along with a control group of randomly selected neurons, setting their activation values to zero. These experiments are performed using the XNLI and Cross-lingual KE (EN (Edit) ALL (Test)) tasks on the BLOOMZ-7b model, with results detailed in Table 2. The most significant decrease in accuracy is observed when the all-shared neurons are deactivated (e.g. 91.6% decrease in the XNLI German testset). While the deactivation of specific and partial-shared neurons also impacts accuracy, these effects are less pronounced compared to those of all-shared neurons. Deactivating neurons randomly has minimal impact on the performance. In order to prove the key role of all-shared neurons across different LLMs, we conduct the ablation experiments with LLAMA2-7b-chat backbone in Appendix A.6.1.
8 Conclusion
In this study, we explored the complex mechanisms of neuron activation within multilingual LLMs, addressing the significant research gap in understanding these models beyond a monolingual context. We developed a fine-grained classification for analysing how neurons respond to different tasks and languages. We categorized neurons into four distinct groups: all-shared, partial-shared, specific, and non-activated. Our research revealed that neuron sharing is heavily influenced by the specific characteristics of tasks and examples, and, surprisingly, it does not always align with language similarity. Furthermore, neurons shared across all languages proved essential for generating accurate responses, highlighting their pivotal role in multilingual processing. Our study improves the understanding of the internal workings of multilingual LLMs and fosters future research in this direction.
9 Limitations
In this paper, we develop a method to analyse neuron behaviors in detail by categorizing them into four distinct neuron types w.r.t the degree of their responses to input languages. Although this enables a fine granularity neuron analysis on LLM backbones across various linguistic characteristics and task complexity, the scope of the experiments can be extended to accommodate larger LLMs with large amounts of parameters (i.e., BLOOMZ-176b) on a more comprehensive range of tasks. While this study demonstrates that the number of languages slightly impacts the percentage of all-shared neurons, it is limited to 10 languages. Exploring the effects of incorporating a larger number of languages into the proposed method warrants further investigation. Moreover, while all-shared neurons in FFN blocks are identified to be of great importance, the specific mechanisms by which they operate remain unclear. Additionally, other network components, for example, attention heads, are not in the scope of this analysis.
Appendix A Appendix
A.1 Detailed Interpretation of Projection in Vocabulary Space
There is a residual connection in the each layer of transformer, where the hidden state is:
In order to analyze the attribution of neurons, we explore how the output distribution in the vocabulary space changes when the representation (before the FFN update) is added with the output of neurons . With the embedding matrix , we map each vector into the vocabulary space . For each token , the probability is calculate with the softmax function:
where is the embedding of , and is the constant softmax normalization factor. The can be viewed as a static score of that is independent of the input to the model. Thus, the projection induces a ranking over the vocabulary. So we use the projection as effective score to detect the responsibility of neurons.
A.2 Tasks
XNLI. Natural Language Inference (Conneau et al., 2018) is a multilingual natural languages inference dataset, containing 5000 items. Each test sample consists of a premise and a hypothesis, requiring an LLM to determine whether a hypothesis is entailed, contradicted, or neutral conditioned on the premise.
Fact Probing. LLMs are used to predict factual answers in response to corresponding probing prompts. A multilingual factual knowledge dataset (mParaRel (Fierro and Søgaard, 2022)) capturing 38 binary relations (e.g., X born-in Y) is used in the analysis. We seletc the relation of “capital” subset (X capital Y) as testset, including 348 items.
Cross-lingual Knowledge Editing (KE). MzsRE (Wang et al., 2023) is a multilingual question-answering dataset, containing 743 items for each language. It provides counterfactual edited knowledge in the context and requires an LLM to produce the corresponding answer according to the context. We evaluate LLMs in two Cross-lingual KE scenarios: 1) EN (Edit) ALL (Test): edit in English and test in other languages and 2) ALL (Edit) EN (Test): edit in other languages and test in English.
A.3 Prompts
For the Fact Probing task, we use the P36 sub-testset, which describe facts of entities in a relation of “capital”. The prompt is framed as “ The capital of {X} is ” where “{X}” is the subject (sovereign state) and LLMs are required to predict the object (capital city). We keep at least three paraphrase prompts from mParaRel for each language to ensure a level of diversity.
For the Natural Language Inference (XNLI) task, we frame the prompt as “ Take the following as truth: {premise} Then the following statement: ‘{hypothesis}’ is ‘true’, ‘false’, or ‘inconclusive’? ”
For the Cross-lingual KE task, we format the prompt as “ {context} Question: {question} Answer: ”. The same language is used for the questions and the answers, but the context is in a different language.
A.4 Supplemental Results of Neuron Behavior Analysis
A.4.1 Neurons Activation Patterns across LLMs
Do the above-mentioned neuron activation patterns change over different LLMs? We further study the neuron activation patterns in another decoder-only multilingual LLM (XGLM) and an encoder-decoder multilingual LLM mT0. The results of XGLM backbones are captured in Figure 12, and the results of mT0-encoder and mT0-decoder are shown in Figure 13 and Figure 14.
The number of active neurons increase in the lower layers, followed by a decrease moving onwards and a rise in the upper layers, despite the absolution values are different from those obtained for BLOOMZ-7b. Moreover, it could be observed that encoder in mT0 involves more all-shared neurons compared to the proportion in decoder.
A.4.2 Neurons Activation Across Examples for Other Tasks
We conduct additional experiments on the XNLI task and Cross-lingual KE (EN (edit) ALL (Test)) task to investigate the type-repeating for neurons. The results shown in Figures 15- 16 are consistent with the results of the Fact Probing task, where only a few neurons retain the same classified types across all examples in the test set.
A.4.3 Neuron Activation Across Languages on XNLI Task
We analyze the shared proportion of German with other languages in partial-shared neurons and the specific neuron ratios for each language derived from the XNLI task in Figure 17. The shared ratio of German with Russian (in different language family) is higher than the ratio of German with French (in the same language family), confirming the conclusion in Section 5.3.
A.4.4 Influence of Model Scale
We investigate neuron activation patterns across the BLOOMZ series with 0.56b, 1b, 3b, 7.1b parameters in a XNLI task. As shown in the results captured in Figure 18, no identifiable pattern difference can be observed to indicate a scale law effect. However, the scale of the model is limited, potentially leading to unreliable results in this experiment. More non-activated neurons in the upper layers of BLOOMZ-7.1b may reflect on a higher level of sparsity for a larger LLM (consistent with Voita et al. (2023); Li et al. (2023)).
A.4.5 Neuron Activation Patterns in Few-shot In-context Learning
According to Wang et al. (2023), in-context learning (ICL) can improve the performance of an LLM under the guidance of few-shot examples in a Cross-lingual KE task. We further explore the impact of few-shot examples on neuron activation patterns. We compare the results of an LLM with 0-shot, 2-shot, 4-shot, 6-shot examples in a Cross-lingual KE (EN (edit) ALL (Test)) task. Four types of neurons in scope have almost identical activation patterns across various few-shot examples (Figure 19). Although in-context examples lead to no observable neuron activation pattern changes, more examples lead to better performances. Could ICL lead to a better neuron activation composition instead of invoking more neurons? We leave this to a future study.
A.4.6 Activation Value across Layers
According to Eq. 1, neurons with larger activation values tend to contribute more to the output. By visualizing neuron activation for a series of thresholds , we can scrutinize the relative importance of various types of activated neurons. The percentage of various types of activated neurons for each activation threshold in the XNLI task are shown in Figure 20. When applying a threshold (>0), there are fewer all-shared, partial-shared, and specific neurons left in the lower layers (layers 0-10) compared to using a lower threshold (i.e., = 0). Under the same threshold scenario, more activated neurons appear in the second half of the model (layers 15-30). It is worth noting these neurons have a higher activation value than neurons activated from a lower threshold (i.e., = 0). Considering both the percentage of activated neurons and their corresponding activation value, it becomes apparent that the neurons in the upper layers contribute more to the output performance.
A.5 Supplemental Results of Neuron Attribution
A.5.1 Contribution Score of Different Tasks
The contribution score of the four types of neurons evaluated on the XNLI task, Cross-lingual KE task, Fact Probing tasks are shown in Figures 24 - 22.
A.5.2 Effective Score of Different Tasks
| all-shared | partial-shared | specific | non-activated | |||||||||
| max | min | mean | max | min | mean | max | min | mean | max | min | mean | |
| en | 1.85 | -0.94 | 0.07 | 0.22 | -0.16 | 0.00 | 0.02 | -0.02 | 0.00 | 0.04 | -0.03 | 0.00 |
| de | 1.03 | -0.60 | 0.02 | 0.13 | -0.13 | 0.00 | 0.07 | -0.03 | 0.00 | 0.02 | -0.01 | 0.00 |
| es | 1.15 | -0.84 | 0.02 | 0.12 | -0.11 | 0.00 | 0.01 | -0.01 | 0.00 | 0.02 | -0.02 | 0.00 |
| fr | 1.06 | -0.78 | 0.01 | 0.15 | -0.11 | 0.00 | 0.03 | -0.04 | 0.00 | 0.02 | -0.02 | 0.00 |
| ru | 0.70 | -0.45 | 0.00 | 0.24 | -0.13 | 0.00 | 0.08 | -0.03 | 0.00 | 0.01 | -0.01 | 0.00 |
| th | 0.50 | -0.90 | 0.00 | 0.17 | -0.10 | 0.00 | 0.03 | -0.05 | 0.00 | 0.01 | -0.01 | 0.00 |
| tr | 0.82 | -0.51 | 0.03 | 0.12 | -0.12 | 0.00 | 0.04 | -0.03 | 0.00 | 0.02 | -0.02 | 0.00 |
| vi | 0.86 | -0.68 | 0.01 | 0.15 | -0.11 | 0.00 | 0.04 | -0.04 | 0.00 | 0.02 | -0.02 | 0.00 |
| zh | 0.52 | -0.42 | 0.00 | 0.17 | -0.20 | 0.00 | 0.08 | -0.07 | 0.00 | 0.02 | -0.01 | 0.00 |
| pt | 1.14 | -0.83 | 0.02 | 0.11 | -0.15 | 0.00 | 0.02 | -0.02 | 0.00 | 0.02 | -0.02 | 0.00 |
The maximum, minimum, and average effective scores of four types of neurons in 10 languages evaluated on the Cross-lingual KE (EN (edit) ALL (Test)) task are shown in Table 3. The maximum effective of four types of neurons across layers evaluated on the Fact Probing task, XNLI task, Cross-lingual KE (ALL (edit) EN (Test)) task are shown in Figures 25 - 27.
A.6 Supplemental Results of Verification on Downstream Tasks
A.6.1 Effects with LLAMA Backbone
| accuracy | pct. | en | de | es | fr | ru | th | tr | vi | zh |
| baseline | 0% | 59.1 | 47.6 | 50.1 | 47.0 | 49.1 | 41.4 | 40.2 | 51.6 | 46.1 |
| w/o. all-shared | 22.42% | 3.0 | 3.6 | 4.4 | 1.9 | 4.7 | 6.9 | 3.6 | 13.5 | 4.8 |
| w/o. partial-shared | 17.48% | 59.1 | 48.4 | 51.5 | 47.9 | 49.7 | 42.9 | 41.5 | 50.8 | 48.0 |
| w/o. specific | 4.75% | 59.2 | 47.3 | 49.9 | 47.0 | 49.1 | 41.9 | 40.1 | 51.4 | 46.2 |
| w/o. non-activated | 55.35% | 30.5 | 13.8 | 12.0 | 11.9 | 12.4 | 5.0 | 14.2 | 13.4 | 5.2 |
| w/o. random | 5% | 58.7 | 47.7 | 50.2 | 48.2 | 49.0 | 41.7 | 40.0 | 49.9 | 45.7 |
| w/o. random | 15% | 52.7 | 44.6 | 47.2 | 46.4 | 44.5 | 38.4 | 40.1 | 48.6 | 45.2 |
| w/o. random | 25% | 46.1 | 42.4 | 41.3 | 43.3 | 40.1 | 34.5 | 39.7 | 38.7 | 40.7 |
| w/o. random | 55% | 28.7 | 30.2 | 28.6 | 30.3 | 25.8 | 19.0 | 27.1 | 28.2 | 25.0 |
In order to further prove the importance of all-shared neurons across LLMs, we conduct the experiments with deactivating neurons on the XNLI task with LLAMA2-7b-chat backbone. The results in Table 4 show that there is more significant effect when all-shared neurons are deactivated. It demonstrates that all-shared neurons play a key role in predicting correct answers across LLMs.

### v3

@@
 However, most of this research is conducted within a monolingual setting, primarily focusing on English.
-Few studies attempt to explore the internal workings of LLMs in multilingual settings.
-In this study, we aim to fill the research gap by examining how neuron activation is shared across tasks and languages.
+Few studies have attempted to explore the internal workings of LLMs in multilingual settings.
+In this study, we aim to fill this research gap by examining how neuron activation is shared across tasks and languages.
 We classify neurons into four distinct categories based on their responses to a specific input across different languages: all-shared, partial-shared, specific, and non-activated.
-This categorization is combined with a study of neuron attribution, i.e.
-the importance of a neuron w.r.t an output.
-Our analysis reveals the following insights: (i) the patterns of neuron sharing are significantly affected by the characteristics of tasks and examples; (ii) neuron sharing does not fully correspond with language similarity; (iii) shared neurons play a vital role in generating responses, especially those shared across all languages.
-These findings shed light on the internal workings of multilingual LLMs and pave the way to the future research.
-We will release the code to foster research in this area.
+Building upon this categorisation, we conduct extensive experiments on three tasks across nine languages using several LLMs and present an in-depth analysis in this work.
+Our findings reveal that: (i) deactivating the all-shared neurons significantly decreases performance; (ii) the shared neurons play a vital role in generating responses, especially for the all-shared neurons; (iii) neuron activation patterns are highly sensitive and vary across tasks, LLMs, and languages.
+These findings shed light on the internal workings of multilingual LLMs and pave the way for future research.
+We release the code to foster research in this area.
+111https://github.com/weixuan-wang123/multilingual-neurons
 Sharing Matters: Analysing Neurons Across Languages and Tasks in LLMs
@@
 1 Introduction
-Large language models (LLMs) have demonstrated remarkable capabilities in recent studies, excelling in both understanding and generating text across various languages (OpenAI, 2023; Zhang et al., 2023; Zhao et al., 2024).
+Large language models (LLMs) have demonstrated remarkable capabilities in recent studies, excelling in both understanding and generating text across various languages (OpenAI, 2023; Zhang et al., 2023; Zhao et al., 2024a).
 Despite their proven effectiveness, the intricate mechanisms underlying their processing remain largely opaque.
@@
 While these studies have analyzed neuron behaviors based on activation states in monolingual settings, there remains a significant gap in our understanding of how neurons behave in multilingual contexts.
-Significant discrepancies have been observed in neuron activation patterns across languages, despite identical semantic outputs, as highlighted in Figure 1.
-This phenomenon raises critical questions about how LLMs manage multilingual representations at the neuronal level.
-In a similar line of reasoning, Bhattacharya and Bojar (2023) categorized neurons into two coarse-grained groups: language-agnostic (shared across languages) and language-specific (unique to a language).
+To address this research gap, recent research attempts to unveil the mechanistic interpretability of multilingual LLMs.
+Bhattacharya and Bojar (2023) categorized neurons into two coarse-grained groups: language-agnostic (shared across languages) and language-specific (unique to a language).
 However this categorization oversimplifies the complexity observed in cross-lingual studies, where neuron overlap varies significantly between languages (Stanczak et al., 2022; Zhao et al., 2023; Liu et al., 2024).
@@
 This underscores the need for a more nuanced, fine-grained classification method to enhance our understanding of the multifaceted roles of neurons in multilingual LLMs.
-To bridge this gap, our research introduces a fine-grained classification of neurons, enabling a detailed exploration of their roles across languages and tasks.
-We systematically analyse neuron behaviors across three distinct tasks in ten languages, including reasoning, fact probing, and question answering tasks, using various diverse model backbones: BLOOMZ-7b, LLaMA2-7b-chat, BLOOM-7b, XGLM, and mT0.
-For a specific input and its corresponding translations in 10 languages, we reformulate the activation state of neurons to four distinctive types to represent multilingual neuron roles (see Section 3).
-All-shared neurons are neurons that remain active for all the translations regardless of language.
-Partial-shared neurons are activated only for translations in certain languages.
-Specific neurons are activated exclusively for translations in one language.
-Non-activated neurons are not activated at all for any the input’s translations.
-We then use the ratio of neurons in each type to analyze the activation pattern across various tasks.
-Furthermore, we evaluate the importance of each type of neurons in predicting answers using contribution scores and effective scores (Geva et al., 2022; Voita et al., 2023).
-Finally, we confirm our analysis through experiments on downstream tasks.
-We provide substantial empirical evidence detailing neuron activation patterns and attributions in this study, leading to several significant findings. Here are the main takeaways:
-Neuron activation patterns are considerably task-specific and example-specific.
-We observe the ratio of four types of neurons is related to the task (see Section 5.1), where the neuron activation pattern of XNLI is significantly different from that of Fact Probing.
-We also demonstrate that the neuron types commonly vary with different test examples (see Section 5.2).
-Similar languages do not always lead to high degree of neuron sharing.
-Our empirical results show that the languages from the same language family, such as German and French, do not exhibit higher degree of neuron sharing, compared to the languages from distinct language families, such as German and Chinese (see Section 5.3).
-Shared neurons are crucial in generating responses.
-Both the contribution scores and the effective scores highlight the significance of all-shared neurons (see Section 6).
-Our experiments further reveal that disabling the all-shared neurons can lead to a substantial decrease (up to 91.6%) in model performance on downstream tasks (see Section 7).
+In this work, our research introduces a fine-grained classification of neurons, enabling a detailed exploration of their functions across languages.
+For a specific English example and its translations in eight other languages, we categorize neurons into four distinctive types (see Figure 1): all-shared neurons, which remain active for all the inputs regardless of language; partial-shared neurons, which are activated only for inputs in certain languages; specific neurons, which are activated exclusively for inputs in one language; and non-activated neurons, which are not activated for any inputs.
+We begin by analysing the importance of each neuron type by deactivating them individually.
+Then we probe their contributions to generating answers using the Generation Impact Score (Geva et al., 2022) and the Correctness Impact Score (Voita et al., 2023).
+Furthermore, by examining the percentage of neurons in each type, we analyse activation patterns to gain insights into the internal workings of LLMs.
+We systematically study neuron behaviours across three distinct tasks, including reasoning, fact probing, and question answering, in nine languages.
+This analysis utilizes diverse model backbones such as BLOOMZ-7b, LLaMA2-7b-chat, BLOOM-7b, and XGLM.
+We provide substantial empirical evidence detailing neuron contributions and activation patterns in this study, leading to several significant findings. Here are the main takeaways:
+All-shared neurons have a significant impact on model performance.
+We individually deactivate each type of neurons in LLMs and observe substantial performance declines (up to 87.39%) across tasks (see Section 5).
+All-shared neurons are crucial in generating responses.
+Both the Generation Impact Score and Correctness Impact Score highlight the significance of the shared neurons in the generation process, and the all-shared neurons make substantially more contributions compared to other neuron types (see Section 6).
+Neuron activation patterns vary across tasks, LLMs, and languages.
+We observe that the patterns of four types of neurons vary across tasks (see Section 7.2) and LLMs (see Section 7.3).
+Moreover, our empirical results show that languages from the same language family do not always exhibit a higher degree of neuron sharing compared with languages from distinct language families (see Section 7.4).
 2 Related Work
@@
 (2024); Kojima et al.
-(2024) classified neurons in an FFN block to language-specific and language-agnostic based on predefined threshold, assuming that a neuron’s classified type remained consistent across different examples.
+(2024) classified neurons in an FFN block to language-specific and language-agnostic based on predefined threshold.
 However, the broad classification into two groups is inadequate for detailed multilingual analysis.
@@
 We investigate neurons’ behaviors across multiple languages and tasks to this end.
-3 Neuron Classification
+3 Fine-Grained Neuron Classification
+| XNLI | KE (EN ALL) | KE (ALL EN) | Fact Probing | |||||||||
+| pct. | pct. | pct. | pct. | |||||||||
+| baseline | 0.00% | 41.99 | 0.00% | 0.00% | 38.39 | 0.00% | 0.00% | 41.74 | 0.00% | 0.00% | 41.98 | 0.00% |
+| w/o. all | 9.92% | 9.38 | -77.66% | 8.71% | 4.84 | -87.39% | 10.17% | 13.19 | -68.40% | 0.28% | 21.86 | -50.31% |
+| w/o. partial | 10.33% | 42.65 | 1.57% | 13.36% | 40.67 | 5.94% | 10.55% | 39.59 | -5.15% | 36.73% | 26.86 | -36.02% |
+| w/o. specific | 3.14% | 42.07 | 0.19% | 4.91% | 40.78 | 6.23% | 3.82% | 40.77 | -2.32% | 16.56% | 12.68 | -67.41% |
+| w/o. non-act. | 76.61% | 35.90 | -14.50% | 73.22% | 21.96 | -42.80% | 75.46% | 19.58 | -53.09% | 46.43% | 26.68 | -36.45% |
+| w/o. random | 5.00% | 42.30 | 0.74% | 5.00% | 30.98 | -19.30% | 5.00% | 41.29 | -1.08% | 1.00% | 37.86 | -9.81% |
+| 15.00% | 43.13 | 2.71% | 15.00% | 31.74 | -17.32% | 15.00% | 42.14 | 0.96% | 15.00% | 35.38 | -15.72% | |
+| 25.00% | 43.98 | 4.74% | 25.00% | 32.40 | -15.60% | 25.00% | 42.28 | 1.29% | 35.00% | 41.78 | -0.48% | |
+| 75.00% | 36.58 | -12.88% | 75.00% | 13.29 | -65.38% | 75.00% | 16.50 | -60.47% | 45.00% | 17.06 | -59.36% | |
 In this section, we provide a detailed description of the 4-way neuron classification that we propose.
 We begin with some background concerning neurons in the FFN block (Section 3.1).
-Following this, we define the four types of neurons (Section 3.2), and explain how we collect activation patterns by averaging the ratio of each neuron type across the entire test set (Section 3.3).
+Following this, we define the four types of neurons (Section 3.2).
 3.1 Neurons in FFN Blocks
 A neuron inside the FFNs is defined as a linear transformation of an input representation followed by a non-linear activation (Tang et al., 2024).
-Every feed-forward block at layer involves two linear transformations separated by a point-wise activation function.
-Biases are omitted for the sake of clarity:
+Every FFN block at layer involves two linear transformations separated by a point-wise activation function.
+Biases are omitted for brevity:
 where , are linear parameter matrices, and is a non-linear activation function, where rows in and columns in are viewed as -dimensional keys and values , respectively.
-We define the behavior of a neuron to be the output of the function immediately after the elementwise nonlinearity.
 is the count of neurons.
@@
 For the -th neuron and corresponding key , value and activation value , we can express this relationship using the following formulation:
-When such a neuron is activated , so it updates the residual stream by pulling out the corresponding value . If a neuron is inactive, the activation value is .
-3.2 Definition of Four Types of Neurons
+Following Voita et al.
+(2023); Bhattacharya and Bojar (2023); Tang et al.
+(2024), we define a neuron as activated when its activation value satisfies .
+Conversely, if the activation value is , the neuron is considered deactivated.
+3.2 Definitions of Four Types of Neurons
 In this work, we categorize the neurons into four types based on their activation values and detail the neuron classification in this section.
@@
 where is the set of all the neurons at the -th layer and is the -th neuron in .
-Furthermore, the non-activated neurons is the set of neurons whose activation value less than or equal to zero in all languages, as follows:
+Furthermore, the non-activated neurons is the set of neurons whose activation value is less than or equal to zero in all languages, as follows:
 Moreover, the specific neurons are the neurons only activated in one specific language and not activated in any other languages, defined as follows:
 Lastly, the remaining neurons are partial-shared neurons as they are activated by inputs from a subset of languages:
-Note that, unlike Bhattacharya and Bojar (2023); Tang et al.
-(2024) who consider the sub-word activation states and thus capture incomplete semantics, we only examine the activation state of the last token, as that is when the LLM performs the prediction task.
-3.3 Collecting Neuron Activation Patterns
-In this section, we explain how to quantify neuron activation patterns.
-Specifically, we measure the percentage of each type of neuron relative to the total number of neurons.
-Given the -th test instance, the percentage of each neuron type at the -th layer can be defined as follows:
-where denotes the number of elements in the set. Consequently, the aggregated neuron activation pattern for one dataset contain instances at the -th layer can be defined as:
+Note that, we only examine the activation state of the last token of the input, as that is when the LLM performs the prediction task.
 4 Experimental Setting
@@
 We perform analysis on neurons in FFN blocks of various LLMs, harnessing their multilingual capabilities in three diverse tasks which consist of multilingual parallel sentences, including XNLI (Conneau et al., 2018), Fact Probing (Fierro and Søgaard, 2022), and Cross-lingual Knowledge Editing (KE) (Wang et al., 2023).
-For the Cross-lingual KE, we analyze the LLMs in two setups, including EN (Edit) ALL (Test) and ALL (Edit) EN (Test).
-More details are described in Appendix A.2.
-These tasks cover 10 diverse languages, including English (en), German (de), Spanish (es), French (fr), Portuguese (pt), Russian (ru), Thai (th), Turkish (tr), Vietnamese (vi), and Chinese (zh).
-Prompts are detailed in Appendix A.3.
-4.2 Base LLMs
-We mainly analyze the behaviors and attributions of neurons in an instruction-finetuned model BLOOMZ-7b (Muennighoff et al., 2023).
-We also include the analysis of other decoder-only models: BLOOM-7b (Scao et al., 2022), LLAMA2-7b-chat (Touvron et al., 2023), XGLM (Lin et al., 2022), and an encoder-decoder model mT0 (Muennighoff et al., 2023) in the Appendix A.4.1.
-5 Behaviors of Four Types of Neurons
-In this section, we present our in-depth analysis with regard to the neuron activation across task (Section 5.1), examples within the same task (Section 5.2), and language (Section 5.3).
-For each task we use parallel test texts in ten languages as input, and record the type of each neuron.
-Subsequently, we calculate and compare the percentage of the four types of neurons compared to the total neurons.
-We use BLOOMZ-7b as the backbone LLM to investigate the behaviors of four types of neurons, and the results of three tasks are shown in Figure 3 respectively.
-5.1 Neuron Activation Across Tasks
-The neuron activation pattern is strongly related to the task performed.
-As shown in Figure 3, non-activated neurons are more prevalent overall than other types.
-However, the pattern observed in the Fact Probing task differs significantly from the other two tasks.
-In the Fact Probing task, there are more partial-shared neurons and specific neurons, whereas the other tasks involve far more all-shared neurons.
-In LABEL:bloom-llama-3tasks, the activation patterns of the BLOOM-7b and LLaMA2-7b-chat LLMs are presented, highlighting variability across models while affirming the consistent influence of the task on neuron activation patterns.
-Additionally, a comparison between BLOOM and BLOOMZ LLMs indicates that instruction fine-tuning in BLOOMZ appears to foster a higher proportion of all-shared neurons, suggesting better alignment across multilingual representations.
-5.2 Neuron Activation Across Examples
-The type of a neuron varies across examples within the same task.
-As shown in Figure 3, we have analysed the percentage of each type of neurons across tasks.
-In this experiment, we now investigate how neurons behave across all examples within one task, aiming to determine whether neurons categorized as a specific type retain their classification across different semantics.
-Take the Fact Probing task as an example, the results are shown in Figure 5 and in Appendix A.4.2 (for other tasks).
-To our surprise, almost no neuron (identified by its index) retains its type (in our four-ways classification) across all examples.
-This finding contradicts prior research that suggested neuronal classification remains unchanged amidst variation in input semantics.
-Our study thus underscores that the classified type of a neuron is sensitive to the semantics of the encountered inputs, even within the same task.
-Non-activated neurons do not stay inactive across examples.
-In Figure 5 we observe that less than 10% of non-activated neurons remain inactive across all examples.
-This small proportion of persistently inactive neurons reflects the distributed nature of knowledge representation in LLMs.
-That is why we should execute caution when pruning neurons, as we may damage the overall performance of an LLM.
-Non-activated neurons do not stay inactive across all tokens in an input.
-Expanding our focus, we examine whether non-activated neurons remain inactive across all tokens of an input, differing from our prior analysis which concentrated only on the last token.
-Specifically, for each input sequence with tokens , we record non-activated neurons of each token, w.r.t the intersection of index.
-A neuron is counted when it stays inactive for every token of one test input sequence.
-As shown in Figure 6, fewer than 0.8% non-activated neurons consistently stay inactive for each token of the input, which further highlights the variability in neuron activity.
-5.3 Neuron Activation Across Languages
-Neuron sharing does not completely align with language similarity.
-In exploring the relationship between language similarity and neuron sharing, we specifically analyse the proportion of shared neurons between German and other languages within the category of partial-shared neurons.
-Additionally, we examine the ratio of specific neurons for each language.
-These analyses are conducted within the Fact Probing task.
-As shown in Figure 7, similar languages do not share more neurons; for instance, German and Chinese (in different language families) have a similar ratio of shared neurons to that of German and French (in the same language family) in the upper layers.
-Furthermore, we observe that there is no consistent pattern with regard to the ratio of specific neurons across the languages studied.
-We leave the further exploration on this matter to the future work.
-More results on XNLI task can be found in Appendix A.4.3.
-More languages encourage greater neuron sharing.
-We also investigate the impact of the number of languages on the percentage of four types of neurons.
-As illustrated in Figure 8, our study include configurations of 3, 5, 7, and 9 languages on the XNLI task.
-Results show a noticeable escalation in the percentage of partial-shared neurons as more languages are introduced, indicating that language configuration affects the results of model’s neuron sharing.
-This analysis emphasizes the potential influence of language selection on multilingual interpretability analysis.
-6 Neuron Attributions
-In the preceding experiments, we examined the proportions of the four types of neurons across tasks and languages.
-In this section, we evaluate importance of neurons in answer generation using contribution scores (Section 6.2), and assess the influence of neurons in predicting correct answers through effective scores (Section 6.3) with BLOOMZ-7b backbone LLM.
-We start with the fundamentals of the contribution analysis method (Section 6.1).
-6.1 Contribution Score and Effective Score
-In this section, we introduce two measures of quantifying the importance of neurons: Contribution Score and Effective Score.
-Contribution Score
-Inspired by Geva et al.
-(2022), in order to judge the importance of neurons in generating answers, we analyze their contributions to the output.
-The contribution score of the -th neuron at -th layer to an FFN output is:
-which is the proportion of its weight to the sum of weights of all neurons in the FFN block. is the absolute value of activation value and is the L2-norm of value .
-Effective Score
-Following Geva et al. (2022) and Voita et al. (2023), the effective score of the -th neuron at -th layer can be defined as:
-where is the embedding of the correct answer .
-A larger has a higher probability to produce the correct answer , while a negative reduces the probability in generating .
-In this way, we can quantify the effect of a neuron on the reference answer.
-We give detailed descriptions about the neuron projection in Appendix A.1.
-Comparison
-The contribution score quantifies how much neurons contribute to the overall generation process, regardless of the correctness of the output, while the effective score incorporates the correct answer’s embedding to measure the neurons’ impact specifically in generating accurate responses.
-Although both metrics assess neuronal influence, the key difference lies in their consideration of answer correctness: the contribution score focuses solely on the generation process, whereas the effective score takes into account the neurons’ role in the accuracy of the resulting answer.
-6.2 The Contribution of Neuron Types
-In this section, we explore the influence of each neuron type using the contribution score described in Section 6.1.
-We examine the proportion of the four types of neurons among the top 5% contribution score under inputs in each language.
-The proportions of neurons in the Cross-lingual KE task are depicted in Figure 9 (and the overall results are shown in Figure 21 in Appendix A.5.1).
-All-shared neurons are the top contributing neurons to outputs.
-Figure 9 depicts a stacked column chart where each segment color represents a specific neuron type’s contribution, collectively totaling 100%.
-It can be observed that all-shared neurons constitute the largest segment, highlighting their significant role irrespective of language input.
-Partial-shared neurons are the second most influential, affecting mainly the latter half of the model.
-It is not surprising that the specific neurons group has limited influence on Cross-lingual KE (EN (Edit) ALL (Test)) outputs as they feature in a particular language type of inputs.
-We further analyze the average and total contribution scores of all neurons across four categories.
-As shown in Figure LABEL:bloomz-contribution_absolute-allspecific_twolangs, all-shared neurons score the highest averages, especially in upper layers, but they constitute only a minor portion of the neuron population (in Figure 3).
-Meanwhile, non-activated and partial-shared neurons contribute significantly to the total scores in the XNLI and fact-probing tasks respectively, predominantly due to their larger numbers.
-The disparity in total contributions across tasks confirms our earlier findings in Section 5.1.
-We leave further exploration regarding to task-specific neuron activation to the future work.
-| language | neuron type | max | min | avg | var |
-| en | all-shared | 1.85 | -0.94 | 0.07 | 0.36 |
-| partial-shared | 0.22 | -0.16 | 1.2e-4 | 1.9e-4 | |
-| specific | 0.02 | -0.02 | 2.5e-4 | 3.5e-5 | |
-| non-activated | 0.04 | -0.03 | 2.1e-4 | 5.8e-6 | |
-| de | all-shared | 1.03 | -0.60 | 0.02 | 0.07 |
-| partial-shared | 0.13 | -0.13 | 6.7e-5 | 7.1e-5 | |
-| specific | 0.07 | -0.03 | 7.5e-5 | 2.3e-5 | |
-| non-activated | 0.02 | -0.01 | 7.0e-5 | 2.9e-6 |
-6.3 The Effectiveness of Neuron Types
-In this section, we explore the effectiveness of each neuron type using the effective score described in Section 6.1.
-All-shared neurons have a greater influence on generating correct answers.
-In the Cross-lingual KE (EN (Edit) ALL (Test)) task, we analyze the effectiveness scores (maximum, minimum, average, and variance) for each neuron type throughout all layers of the BLOOMZ model, as shown in Table 1.
-All-shared neurons display both the highest maximum and the lowest minimum scores, suggesting they are pushed (or eliminated) strongly by the activation functions.
-By comparison, while all-shared neurons and partial-shared neurons show a wide variance in scores (1.85 vs.
--0.94 and 0.22 vs.
--0.16, respectively, in English), specific neurons and non-activated neurons have much tighter score ranges ( 0.07).
-Additionally, we demonstrate the maximum effective scores across layers in Figure 11.
-It could be observed the neurons in the upper layers play a more crucial role in determining model performance.
+For the Cross-lingual KE, we analyse the LLMs in two setups, including EN (Edit) ALL (Test) and ALL (Edit) EN (Test).
+These test sets across languages are translated from the original English test set.
+More details are described in Appendix C.
+These tasks cover nine diverse languages, including English (en), German (de), Spanish (es), French (fr), Russian (ru), Thai (th), Turkish (tr), Vietnamese (vi), and Chinese (zh).
+Prompts are detailed in Appendix D.
+4.2 Model Backbones
+We mainly analyse the contributions and activation patterns of neurons in an instruction-finetuned multilingual model BLOOMZ-7b (Muennighoff et al., 2023).
+We also include the analysis of other multilingual LLMs: BLOOM-7b (Scao et al., 2022), LLAMA2-7b-chat (Touvron et al., 2023), and XGLM (Lin et al., 2022).
+We use one NVIDIA A100 (40G) for all experiments.
+5 Shared Neurons Are Crucial to Performance
 | settings | pct. | en | de | es | fr | ru | th | tr | vi | zh |
@@
 | baseline | 0% | 53.8 | 41.8 | 50.3 | 49.0 | 47.6 | 40.9 | 34.9 | 50.5 | 51.1 |
-| w/o. all-shared | 9.85% | 16.7 | 3.5 | 10.1 | 10.0 | 6.6 | 9.0 | 1.4 | 12.1 | 14.5 |
-| w/o. partial-shared | 10.30% | 52.9 | 40.4 | 49.7 | 47.6 | 49.2 | 40.3 | 36.1 | 50.0 | 50.0 |
-| w/o. specific | 3.12% | 53.7 | 41.7 | 50.3 | 48.9 | 47.4 | 40.6 | 35.3 | 50.4 | 49.3 |
-| w/o. non-activated | 76.73% | 36.6 | 31.6 | 33.6 | 33.4 | 29.5 | 31.3 | 28.3 | 34.5 | 23.5 |
+| w/o. all | 9.9% | 16.7 | 3.5 | 10.1 | 10.0 | 6.6 | 9.0 | 1.4 | 12.1 | 14.5 |
+| w/o. partial | 10.3% | 52.9 | 40.4 | 49.7 | 47.6 | 49.2 | 40.3 | 36.1 | 50.0 | 50.0 |
+| w/o. specific | 3.1% | 53.7 | 41.7 | 50.3 | 48.9 | 47.4 | 40.6 | 35.3 | 50.4 | 49.3 |
+| w/o. non-act. | 76.7% | 36.6 | 31.6 | 33.6 | 33.4 | 29.5 | 31.3 | 28.3 | 34.5 | 23.5 |
 | w/o. random | 5% | 53.2 | 42.2 | 50.7 | 48.8 | 47.4 | 40.2 | 34.5 | 50.1 | 50.9 |
-| w/o. random | 15% | 53.1 | 41.8 | 50.1 | 48.9 | 47.3 | 40.8 | 33.8 | 50.1 | 50.4 |
-| w/o. random | 25% | 52.6 | 41.7 | 50.3 | 48.8 | 46.0 | 38.8 | 36.2 | 50.7 | 49.7 |
-| w/o. random | 75% | 36.0 | 28.7 | 40.7 | 36.7 | 28.9 | 25.4 | 23.0 | 38.5 | 32.9 |
+| 15% | 53.1 | 41.8 | 50.1 | 48.9 | 47.3 | 40.8 | 33.8 | 50.1 | 50.4 | |
+| 25% | 52.6 | 41.7 | 50.3 | 48.8 | 46.0 | 38.8 | 36.2 | 50.7 | 49.7 | |
+| 75% | 36.0 | 28.7 | 40.7 | 36.7 | 28.9 | 25.4 | 23.0 | 38.5 | 32.9 | |
 | Cross-lingual KE (EN (Edit) ALL (Test)) task | ||||||||||
-| baseline | 0% | 96.2 | 55.1 | 48.9 | 49.5 | 30.6 | 6.3 | 38.8 | 51.7 | 33.4 |
-| w/o. all-shared | 8.69% | 11.0 | 4.9 | 6.1 | 4.9 | 1.9 | 0.4 | 1.5 | 6.1 | 2.9 |
-| w/o. partial-shared | 13.20% | 90.2 | 51.7 | 48.9 | 48.9 | 25.4 | 5.5 | 35.5 | 50.9 | 38.4 |
-| w/o. specific | 4.87% | 96.1 | 54.4 | 48.7 | 48.9 | 30.4 | 6.3 | 37.9 | 51.7 | 28.5 |
-| w/o. non-activated | 73.20% | 36.1 | 15.8 | 18.2 | 17.8 | 1.9 | 9.8 | 19.9 | 10.6 | 16.3 |
+| baseline | 0% | 96.2 | 48.8 | 36.9 | 49.5 | 24.6 | 6.3 | 38.8 | 49.4 | 33.4 |
+| w/o. all | 8.7% | 11.0 | 4.9 | 6.1 | 4.9 | 1.9 | 0.4 | 1.5 | 6.1 | 2.9 |
+| w/o. partial | 13.3% | 90.2 | 51.7 | 46.9 | 48.9 | 25.4 | 5.5 | 35.5 | 50.9 | 38.4 |
+| w/o. specific | 4.9% | 96.1 | 54.4 | 48.7 | 48.9 | 30.4 | 6.3 | 37.9 | 51.7 | 28.5 |
+| w/o. non-act. | 73.2% | 36.1 | 15.8 | 18.2 | 17.8 | 1.9 | 9.8 | 19.9 | 10.6 | 16.3 |
 | w/o. random | 5% | 96.1 | 46.9 | 36.6 | 40.4 | 0.8 | 4.4 | 28.7 | 40.8 | 10.1 |
-| w/o. random | 15% | 94.8 | 47.1 | 36.1 | 39.4 | 0.8 | 4.3 | 28.4 | 40.4 | 11.1 |
-| w/o. random | 25% | 91.5 | 46.8 | 36.2 | 38.6 | 1.1 | 4.4 | 27.9 | 40.4 | 12.1 |
-| w/o. random | 75% | 11.1 | 5.5 | 9.3 | 11.3 | 0.1 | 2.7 | 1.9 | 8.9 | 7.1 |
-7 Verification on Downstream Tasks
-We analyse the behaviors of different neuron types in Section 5 and explore the contribution and effectiveness of various types of neurons in Section 6, demonstrating that neuron shared across languages and tasks are of vital importance from an internal representation perspective.
-In this section, we conduct experiments on downstream tasks to verify our analytical findings regarding the significance of neuron sharing.
-We assess accuracy changes by selectively deactivating four specific types of neurons along with a control group of randomly selected neurons, setting their activation values to zero.
-These experiments are performed using the XNLI and Cross-lingual KE (EN (Edit) ALL (Test)) tasks on the BLOOMZ-7b model, with results detailed in Table 2.
-The most significant decrease in accuracy is observed when the all-shared neurons are deactivated (e.g.
-91.6% decrease in the XNLI German testset).
-While the deactivation of specific and partial-shared neurons also impacts accuracy, these effects are less pronounced compared to those of all-shared neurons.
-Deactivating neurons randomly has minimal impact on the performance.
-In order to prove the key role of all-shared neurons across different LLMs, we conduct the ablation experiments with LLAMA2-7b-chat backbone in Appendix A.6.1.
+| 15% | 94.8 | 47.1 | 36.1 | 39.4 | 0.8 | 4.3 | 28.4 | 40.4 | 11.1 | |
+| 25% | 91.5 | 46.8 | 36.2 | 38.6 | 1.1 | 4.4 | 27.9 | 40.4 | 12.1 | |
+| 75% | 11.1 | 5.5 | 9.3 | 11.3 | 0.1 | 2.7 | 1.9 | 8.9 | 7.1 | |
+| Cross-lingual KE (ALL (Edit) EN (Test)) task | ||||||||||
+| baseline | 0% | 96.2 | 55.1 | 49.2 | 49.5 | 30.6 | 9.2 | 39.3 | 51.7 | 36.6 |
+| w/o. all | 10.2% | 24.4 | 19.5 | 13.8 | 13.1 | 8.0 | 1.4 | 14.5 | 19.9 | 7.1 |
+| w/o. partial | 10.5% | 85.1 | 51.3 | 47.8 | 48.0 | 25.4 | 5.2 | 35.1 | 51.4 | 36.1 |
+| w/o. specific | 3.8% | 96.1 | 54.4 | 48.7 | 48.9 | 29.7 | 6.3 | 38.0 | 51.8 | 30.0 |
+| w/o. non-act. | 75.5% | 19.2 | 19.8 | 15.5 | 14.7 | 11.4 | 2.0 | 18.2 | 12.0 | 7.5 |
+| w/o. random | 5% | 95.6 | 54.2 | 48.7 | 50.2 | 29.9 | 6.5 | 38.5 | 51.3 | 33.0 |
+| 15% | 93.9 | 56.1 | 49.1 | 49.9 | 29.2 | 6.4 | 38.2 | 51.0 | 32.6 | |
+| 25% | 91.3 | 55.9 | 48.5 | 49.3 | 28.3 | 6.8 | 37.7 | 50.3 | 29.7 | |
+| 75% | 10.2 | 17.0 | 11.7 | 15.7 | 4.7 | 1.1 | 7.8 | 14.8 | 7.0 | |
+| Fact Probing task | ||||||||||
+| baseline | 0% | 72.4 | 41.6 | 56.6 | 58.1 | 37.3 | 5.7 | 39.3 | 57.4 | 51.4 |
+| w/o. all | 0.2% | 43.4 | 12.9 | 34.4 | 22.4 | 11.4 | 5.2 | 15.2 | 34.5 | 29.0 |
+| w/o. partial | 36.7% | 43.3 | 20.8 | 31.2 | 30.9 | 14.4 | 2.8 | 24.5 | 34.6 | 29.4 |
+| w/o. specific | 16.6% | 18.1 | 9.1 | 30.1 | 7.7 | 5.7 | 5.7 | 9.4 | 12.3 | 22.1 |
+| w/o. non-act. | 46.4% | 42.5 | 27.6 | 39.9 | 28.1 | 1.7 | 0.0 | 22.9 | 40.2 | 17.5 |
+| w/o. random | 1% | 76.4 | 50.6 | 48.6 | 56.0 | 3.2 | 0.0 | 36.2 | 59.5 | 47.1 |
+| 15% | 71.5 | 48.5 | 45.6 | 63.5 | 4.3 | 0.0 | 22.7 | 44.5 | 38.2 | |
+| 35% | 77.3 | 51.4 | 50.3 | 56.3 | 4.6 | 0.0 | 37.1 | 57.5 | 48.3 | |
+| 45% | 29.0 | 21.3 | 16.4 | 17.2 | 0.3 | 0.0 | 9.5 | 25.9 | 6.0 | |
+In this section, we explore how different neuron types affect the performance of the BLOOMZ-7b model by selectively deactivating specific groups of neurons.
+By setting the activation values of these neurons to zero, we assess their impact on the model’s output across various tasks.
+Specifically, we compare the effects of deactivating four distinct types of neurons and include a control group of randomly selected neurons to evaluate their respective contributions to the model performance.
+Our experiments involve tasks such as XNLI, cross-lingual KE, and fact probing.
+All-shared neurons play a crucial role in model performance across different tasks.
+As shown in Table 1, we observe that all-shared neurons significantly contribute to the model’s performance across various tasks.
+For instance, in the Cross-lingual KE (EN (Edit) ALL (Test)) task, deactivating the all-shared neurons, which account for only 8.71% of the total neurons, results in an 87.39% decrease in accuracy.
+Moreover, for the Fact Probing task, deactivating the all-shared neurons, which constitute only 0.28% of the total neurons, causes a substantial 50.31% performance drop.
+Furthermore, deactivating the specific neurons, which account for 16.56% of the total neurons, leads to the largest performance decline of 67.41%.
+In comparison, deactivating a comparable number of random selected neurons typically results in smaller performance drops, suggesting that all-shared neurons are crucial to the performance.
+Deactivating neurons does not always result in performance declines.
+Interestingly, we sometimes observe small performance gains when a small number of neurons are deactivated, as shown in Table 1, regardless of the neuron type.
+To explore this phenomenon further, we provide a breakdown of the results by language in Table 2.
+Our analysis reveals that only deactivating the all-shared neurons consistently leads to a decline in model performance across various tasks and languages.
+In contrast, deactivating either partial-shared neurons or specific neurons can occasionally improve performance for certain languages.
+For example, in the Cross-lingual KE (EN (Edit) ALL (Test)) task, we observe substantial performance improvements in German (de), Spanish (es), and Chinese (zh) when the partial-shared neurons are deactivated.
+We hypothesize that this phenomenon stems from knowledge conflicts encoded in the LLM (Xu et al., 2024).
+By deactivating certain neurons, these knowledge conflicts may be mitigated, resulting in enhanced performance.
+It is important to note that we conduct similar experiments using the LLAMA2-7b-chat and present the results in Appendix E.
+These additional experiments yield observations and conclusions consistent with those using BLOOMZ-7b.
+6 Probing Neuron Contributions
+We demonstrate the significant role of neurons shared across languages, particularly the all-shared neurons, in generating answers, as discussed in Section 5.
+To gain a deeper understanding of the model’s behavior, we conduct further analysis using two metrics: Generation Impact Score and Correctness Impact Score.
+First, we introduce the definitions of these two metrics in Section 6.1.
+Then, we analyse and quantify the contributions of each type of neuron in Section 6.2 and Section 6.3, respectively.
+6.1 Generation Impact Score and Correctness Impact Score
+In this section, we introduce two measures to quantify the contributions of neurons: Generation Impact Score and Correctness Impact Score.
+Generation Impact Score
+Inspired by Geva et al. (2022), the Generation Impact Score () evaluates the importance of neurons in generating answers. For the -th neuron at -th layer, the is defined as:
+which is the proportion of its weight to the sum of weights of all neurons in the FFN block. is the absolute value of activation value and is the L2-norm of value (the column in matrix ).
+Correctness Impact Score
+Following Geva et al. (2022) and Voita et al. (2023), Correctness Impact Score () assesses a neuron’s influence on generating the correct answer.
+where is the embedding of the correct answer .
+A larger has a higher probability to produce the correct answer , while a negative reduces the probability in generating .
+Detailed descriptions of the neuron projection are provided in Appendix B.
+| all-shared | partial-shared | specific | non-activated | |||||||||||||
+| max | min | mean | var | max | min | mean | var | max | min | mean | var | max | min | mean | var | |
+| en | 1.85 | -0.94 | 0.07 | 0.36 | 0.22 | -0.16 | 1.2e-4 | 1.9e-4 | 0.02 | -0.02 | 2.5e-4 | 3.5e-5 | 0.04 | -0.03 | 2.1e-4 | 5.8e-6 |
+| de | 1.03 | -0.60 | 0.02 | 0.07 | 0.13 | -0.13 | 6.7e-5 | 7.1e-5 | 0.07 | -0.03 | 2.3e-5 | 2.3e-5 | 0.02 | -0.01 | 2.9e-6 | 2.9e-6 |
+| es | 1.15 | -0.84 | 0.02 | 0.07 | 0.12 | -0.11 | 1.3e-4 | 6.3e-5 | 0.01 | -0.01 | 9.4e-5 | 7.6e-6 | 0.02 | -0.02 | 7.7e-5 | 3.1e-6 |
+| fr | 1.06 | -0.78 | 0.01 | 0.05 | 0.15 | -0.11 | 2.1e-4 | 7.8e-5 | 0.03 | -0.04 | 3.6e-5 | 9.9e-6 | 0.02 | -0.02 | 5.6e-5 | 2.8e-6 |
+| ru | 0.70 | -0.45 | 3.3e-3 | 5.9e-3 | 0.24 | -0.13 | 2.6e-4 | 7.6e-5 | 0.08 | -0.03 | 1.1e-4 | 1.8e-5 | 0.01 | -0.01 | 3.9e-5 | 1.7e-6 |
+| th | 0.50 | -0.90 | 2.6e-3 | 0.02 | 0.17 | -0.10 | 2.2e-4 | 6.8e-5 | 0.03 | -0.05 | 3.4e-5 | 1.8e-5 | 0.01 | -0.01 | 7.1e-5 | 1.9e-6 |
+| tr | 0.82 | -0.51 | 0.03 | 0.07 | 0.12 | -0.12 | 1.6e-4 | 7.4e-5 | 0.04 | -0.03 | 6.6e-5 | 1.1e-5 | 0.02 | -0.02 | 9.1e-5 | 3.4e-6 |
+| vi | 0.86 | -0.68 | 6.8e-3 | 0.03 | 0.15 | -0.11 | 9.3e-5 | 6.8e-5 | 0.04 | -0.04 | 2.8e-5 | 1.3e-5 | 0.02 | -0.02 | 3.2e-5 | 2.6e-6 |
+| zh | 0.52 | -0.42 | 1.9e-3 | 0.02 | 0.17 | -0.20 | 1.5e-4 | 7.7e-5 | 0.08 | -0.07 | 8.5e-5 | 2.6e-5 | 0.02 | -0.01 | 1.7e-5 | 3.1e-6 |
+Comparison
+While both Generation Impact Score () and Correctness Impact Score () measure neuronal influence, they serve different purposes.
+The quantifies a neuron’s overall contribution to the generation process, regardless of output correctness.
+In contrast specifically measures a neuron’s impact on producing accurate responses by incorporating the correct answer’s embedding.
+Thus, the key distinction lies in their consideration of answer correctness: focuses on general generation ability, whereas emphasizes correctness.
+6.2 The Generation Impact of Neuron Types
+In this section, we explore the contribution of each neuron type using the Generation Impact Score () described in Section 6.1.
+All-shared neurons have the greatest impact on generation outputs.
+As shown in Figure 3(a), we analyse the across layers on the English and German test sets of three tasks (with overall results provided in Appendix F).
+For both English and German, it can be observed that the all-shared neurons almost always achieve the highest across all layers, indicating their significant influence on the model’s output generation.
+The partial-shared neurons are the second most influential, particularly in the upper layers.
+Notably, there is a decrease in the influence of all-shared neurons between layers 5 and 10.
+This can be attributed to the fact that assesses the impact on generating answers, while the lower layers are primarily responsible for input understanding (Zhao et al., 2024b).
+Consequently, all types of neurons exhibit lower in these layers.
+Moreover, previous studies have demonstrated that higher layers capture more abstract, high-level information essential for generation (Gao et al., 2024).
+These findings suggest that shared neurons play a more significant role in the model’s generation capabilities.
+6.3 The Correctness Impact of Neuron Types
+In this section, we assess the effectiveness of each neuron type using the Correctness Impact Score () described in Section 6.1.
+All-shared neurons have the greatest impact on generating correct answers.
+In the Cross-lingual KE (EN (Edit) ALL (Test)) task, we present the maximum, minimum, average, and variance of for each neuron type across all layers of the BLOOMZ-7b model, as shown in Table 3.
+The results reveal that all-shared neurons have both the highest maximum and the lowest minimum values, indicating that they have strong impact on generating correct outputs.
+While all-shared and partial-shared neurons display a wide variance in (e.g., 1.85 vs.
+-0.94 and 0.22 vs.
+-0.16 in English, respectively), specific neurons and non-activated neurons exhibit much narrower score ranges (approximately 0.07).
+Furthermore, the all-shared neurons also exhibit the largest mean and variance of among all kinds of neurons.
+In conclusion, these findings presented in Section 6.2 and Section 6.3 demonstrate that all-shared neurons also have the greatest impact on generating both answers and correct answers, highlighting their importance in the model’s performance across different languages and tasks.
+7 Understanding Neuron Activations
+We demonstrate in Section 5 that shared neurons have a significant impact on model performance and investigate their influence on the generation process in Section 6.
+However, the inner patterns of neurons across layers remain unexplored.
+In this section, we firstly introduce the measure of quantifying neuron activation in Section 7.1, and then we further illustrate how neuron activation patterns vary across tasks (Section 7.2), LLMs (Section 7.3) and languages (Section 7.4).
+7.1 Measuring Neuron Activation
+In this section, we explain how to quantify neuron activation patterns based on the definitions in Section 3.2.
+Specifically, we measure the percentage of each type of neuron relative to the total number of neurons.
+Given the -th test instance, the percentage of each neuron type at the -th layer can be defined as follows:
+where denotes the number of elements in the set. Consequently, the aggregated neuron activation pattern at the -th layer for one dataset containing instances can be defined as:
+7.2 Neuron Activations Across Tasks
+Neuron activations are task-related.
+As shown in Figure 3, non-activated neurons are typically more prevalent than other types of neurons, except in the Fact Probing task.
+In this task, there are more partial-shared neurons and specific neurons, with a negligible amount of all-shared neurons, whereas other tasks involve far more all-shared neurons.
+Referring to Table 1, deactivating the specific neurons and all-shared neurons results in the largest and second largest performance declines.
+These findings demonstrate that the some factual knowledges in LLMs are language-specific and minimally shared across languages, while others are universally shared.
+We leave more in-depth investigation to the future work.
+Neuron sharing peaks at early layers for universal features, declining later for specific ones.
+We present the percentage of each neuron type at each layer in Figure 3.
+The number of all-shared neurons and partial-shared neurons typically peaks between the 5th and 10th layers and then gradually decreases in subsequent layers.
+This trend can be explained by the functional roles of different layers in the model.
+The initial layers, which are closer to the input data, primarily focus on capturing low-level features such as basic lexical and syntactic patterns.
+As the network progresses to the later layers (between the 5th and 10th layers), it begins to learn abstract concepts that are relatively universal across different tasks and languages.
+This universality leads to a higher number of shared neurons in these layers.
+In contrast, the higher layers specialize in task-specific features and nuances unique to each task, resulting in a decline in neuron sharing.
+These findings highlight the importance of neuron sharing in LLMs, as shared neurons in the early layers facilitate the transfer of universal knowledge across tasks and languages.
+They also align with previous research (Yosinski et al., 2014; de Vries et al., 2020; Zhao et al., 2024b; Gao et al., 2024).
+7.3 Neuron Activations Across LLMs
+Different LLMs exhibit different neuron activation patterns.
+To investigate whether neuron activation patterns vary across different multilingual LLMs, we present additional results from LLaMA2-7b-chat in Figure 4.
+Our analysis reveals that the activation patterns in LLaMA2-7b-chat differ significantly from those observed in BLOOMZ-7b, highlighting the variability across models.
+Notably, LLaMA2-7b-chat demonstrates a higher degree of neuron sharing, particularly for partial-shared neurons.
+This phenomenon can be attributed to the English-centric nature of LLaMA2-7b-chat.
+When processing multilingual inputs, the model heavily relies on knowledge transfer from English to other languages, resulting in a substantial number of partial-shared neurons.
+We also present additional results using XGLM (Lin et al., 2022) in Appendix G, aligning with our observations.
+Instruction finetuned LLMs exhibit larger proportion of the all-shared neurons.
+We conduct additional experiments using the foundation model BLOOM-7b to explore the impact of instruction finetuning on neuron activation patterns.
+As shown in Figure 5, the instruction-finetuned BLOOMZ-7b demonstrates a higher percentage of all-shared neurons compared to BLOOM-7b.
+This observation suggests that instruction finetuning may encourage neuron sharing within LLMs, potentially aligning their internal representations across languages.
+Therefore, instruction-finetuned LLMs, such as BLOOMZ-7b, generally outperform their foundational counterparts.
+7.4 Neuron Activations Across Languages
+Neuron sharing does not completely align with language similarity.
+We investigate the relationship between language similarity and neuron sharing by analysing the proportion of partial-shared neurons for language pairs involving German and several other languages on the Fact Probing task.
+As shown in Figure 6, our findings reveal that similar languages (e.g., German and French) do not always exhibit higher levels of neuron sharing.
+For instance, the proportion of partial-shared neurons between German and Chinese is nearly identical to that between German and French, despite German and French both belonging to the Indo-European language family, while Chinese belongs to the Sino-Tibetan language family.
+Furthermore, we observe no consistent pattern in the percentage of specific neurons across the languages studied, suggesting that neuron specialization may not directly correlate with language similarity.
+We leave further exploration of this phenomenon to future work.
+Additional results on XNLI task are in Appendix I.
+Furthermore, we conduct ablation studies to investigate the impact of two key factors on the neuron activation patterns: the size of the backbone model with 0.56b, 1b, 3b, 7b parameters (Appendix J.).
 8 Conclusion
@@
 We categorized neurons into four distinct groups: all-shared, partial-shared, specific, and non-activated.
-Our research revealed that neuron sharing is heavily influenced by the specific characteristics of tasks and examples, and, surprisingly, it does not always align with language similarity.
-Furthermore, neurons shared across all languages proved essential for generating accurate responses, highlighting their pivotal role in multilingual processing.
+Our research revealed that neurons shared across all languages proved essential for generating accurate responses, highlighting their pivotal role in multilingual processing.
+Furthermore, we demonstrate that neuron sharing is task-related, and, it does not always align with language similarity.
 Our study improves the understanding of the internal workings of multilingual LLMs and fosters future research in this direction.
-9 Limitations
-In this paper, we develop a method to analyse neuron behaviors in detail by categorizing them into four distinct neuron types w.r.t the degree of their responses to input languages.
-Although this enables a fine granularity neuron analysis on LLM backbones across various linguistic characteristics and task complexity, the scope of the experiments can be extended to accommodate larger LLMs with large amounts of parameters (i.e., BLOOMZ-176b) on a more comprehensive range of tasks.
-While this study demonstrates that the number of languages slightly impacts the percentage of all-shared neurons, it is limited to 10 languages.
-Exploring the effects of incorporating a larger number of languages into the proposed method warrants further investigation.
-Moreover, while all-shared neurons in FFN blocks are identified to be of great importance, the specific mechanisms by which they operate remain unclear.
-Additionally, other network components, for example, attention heads, are not in the scope of this analysis.
-Appendix A Appendix
-A.1 Detailed Interpretation of Projection in Vocabulary Space
+Acknowledgement
+This work is funded by EU Horizon Europe (HE) Research and Innovation programme grant No 101070631, and UK Research and Innovation under the UK HE funding grant No 10039436.
+The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/).
+Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
+Appendix A Detailed Interpretation of Projection in Vocabulary Space
 There is a residual connection in the each layer of transformer, where the hidden state is:
@@
 So we use the projection as effective score to detect the responsibility of neurons.
-A.2 Tasks
+Appendix B Tasks
 XNLI.
-Natural Language Inference (Conneau et al., 2018) is a multilingual natural languages inference dataset, containing 5000 items.
+Natural Language Inference is a multilingual natural languages inference dataset, containing 5000 items.
 Each test sample consists of a premise and a hypothesis, requiring an LLM to determine whether a hypothesis is entailed, contradicted, or neutral conditioned on the premise.
@@
 LLMs are used to predict factual answers in response to corresponding probing prompts.
-A multilingual factual knowledge dataset (mParaRel (Fierro and Søgaard, 2022)) capturing 38 binary relations (e.g., X born-in Y) is used in the analysis.
+A multilingual factual knowledge dataset (mParaRel) capturing 38 binary relations (e.g., X born-in Y) is used in the analysis.
 We seletc the relation of “capital” subset (X capital Y) as testset, including 348 items.
 Cross-lingual Knowledge Editing (KE).
-MzsRE (Wang et al., 2023) is a multilingual question-answering dataset, containing 743 items for each language.
+MzsREis a multilingual question-answering dataset, containing 743 items for each language.
 It provides co unterfactual edited knowledge in the context and requires an LLM to produce the corresponding answer according to the context.
 We evaluate LLMs in two Cross-lingual KE scenarios: 1) EN (Edit) ALL (Test): edit in English and test in other languages and 2) ALL (Edit) EN (Test): edit in other languages and test in English.
-A.3 Prompts
+Appendix C Prompts
 For the Fact Probing task, we use the P36 sub-testset, which describe facts of entities in a relation of “capital”.
@@
 The same language is used for the questions and the answers, but the context is in a different language.
-A.4 Supplemental Results of Neuron Behavior Analysis
-A.4.1 Neurons Activation Patterns across LLMs
-Do the above-mentioned neuron activation patterns change over different LLMs?
-We further study the neuron activation patterns in another decoder-only multilingual LLM (XGLM) and an encoder-decoder multilingual LLM mT0.
-The results of XGLM backbones are captured in Figure 12, and the results of mT0-encoder and mT0-decoder are shown in Figure 13 and Figure 14.
-The number of active neurons increase in the lower layers, followed by a decrease moving onwards and a rise in the upper layers, despite the absolution values are different from those obtained for BLOOMZ-7b.
-Moreover, it could be observed that encoder in mT0 involves more all-shared neurons compared to the proportion in decoder.
-A.4.2 Neurons Activation Across Examples for Other Tasks
-We conduct additional experiments on the XNLI task and Cross-lingual KE (EN (edit) ALL (Test)) task to investigate the type-repeating for neurons.
-The results shown in Figures 15- 16 are consistent with the results of the Fact Probing task, where only a few neurons retain the same classified types across all examples in the test set.
-A.4.3 Neuron Activation Across Languages on XNLI Task
+Appendix D Supplemental Results on Deactivating Neurons
+| settings | pct. | en | de | es | fr | ru | th | tr | vi | zh |
+| baseline | 0% | 59.1 | 47.6 | 50.1 | 47.0 | 49.1 | 41.4 | 40.2 | 51.6 | 46.1 |
+| w/o. all-shared | 22.42% | 3.0 | 3.6 | 4.4 | 1.9 | 4.7 | 6.9 | 3.6 | 13.5 | 4.8 |
+| w/o. partial-shared | 17.48% | 59.1 | 48.4 | 51.5 | 47.9 | 49.7 | 42.9 | 41.5 | 50.8 | 48.0 |
+| w/o. specific | 4.75% | 59.2 | 47.3 | 49.9 | 47.0 | 49.1 | 41.9 | 40.1 | 51.4 | 46.2 |
+| w/o. non-activated | 55.35% | 30.5 | 13.8 | 12.0 | 11.9 | 12.4 | 5.0 | 14.2 | 13.4 | 5.2 |
+| w/o. random | 5% | 58.7 | 47.7 | 50.2 | 48.2 | 49.0 | 41.7 | 40.0 | 49.9 | 45.7 |
+| 15% | 52.7 | 44.6 | 47.2 | 46.4 | 44.5 | 38.4 | 40.1 | 48.6 | 45.2 | |
+| 25% | 46.1 | 42.4 | 41.3 | 43.3 | 40.1 | 34.5 | 39.7 | 38.7 | 40.7 | |
+| 55% | 28.7 | 30.2 | 28.6 | 30.3 | 25.8 | 19.0 | 27.1 | 28.2 | 25.0 |
+In order to further prove the importance of all-shared neurons across LLMs, we conduct the experiments with deactivating neurons on the XNLI task with LLAMA2-7b-chat backbone.
+The results in Table 4 show that there is more significant decline when all-shared neurons are deactivated.
+It demonstrates that all-shared neurons play a key role in predicting correct answers across LLMs.
+Appendix E Generation Impact Score of Different Tasks
+The Generation Impact Score of the four types of neurons evaluated on the Cross-lingual KE (EN (edit) ALL (Test)) and XNLI tasks across languages are shown in Figure 7 and Figure 8.
+Appendix F Supplemental Results on Neurons Activation Patterns across LLMs
+We further study the neuron activation patterns in another multilingual LLM (XGLM). The results of XGLM backbone are captured in Figure 9.
+Appendix G Supplemental Results on Neurons Activation Patterns of Foundation LLM BLOOM-7b
+We further explore the neuron activation patterns across various tasks in the foundation LLM (BLOOM-7b). The results of BLOOM-7b backbone are captured in Figure 10.
+Appendix H Neuron Activation Across Languages on XNLI Task
 We analyze the shared proportion of German with other languages in partial-shared neurons and the specific neuron ratios for each language derived from the XNLI task in Figure 11.
 The shared ratio of German with Russian (in different language family) is higher than the ratio of German with French (in the same language family), confirming the conclusion in Section 7.4.
-A.4.4 Influence of Model Scale
-We investigate neuron activation patterns across the BLOOMZ series with 0.56b, 1b, 3b, 7.1b parameters in a XNLI task.
+Appendix I Influence of Model Scale
+We investigate neuron activation patterns across the BLOOMZ series with 0.56b, 1b, 3b, 7b parameters in a XNLI task.
 As shown in the results captured in Figure 12, no identifiable pattern difference can be observed to indicate a scale law effect.
 However, the scale of the model is limited, potentially leading to unreliable results in this experiment.
-More non-activated neurons in the upper layers of BLOOMZ-7.1b may reflect on a higher level of sparsity for a larger LLM (consistent with Voita et al.
-(2023); Li et al.
-(2023)).
-A.4.5 Neuron Activation Patterns in Few-shot In-context Learning
-According to Wang et al.
-(2023), in-context learning (ICL) can improve the performance of an LLM under the guidance of few-shot examples in a Cross-lingual KE task.
+More non-activated neurons in the upper layers of BLOOMZ-7b may reflect on a higher level of sparsity for a larger LLM.
+Appendix J Neuron Activation Patterns in Few-shot In-context Learning
+In-context learning (ICL) can improve the performance of an LLM under the guidance of few-shot examples in a Cross-lingual KE task.
 We further explore the impact of few-shot examples on neuron activation patterns.
@@
 We leave this to a future study.
-A.4.6 Activation Value across Layers
-According to Eq.
-1, neurons with larger activation values tend to contribute more to the output.
-By visualizing neuron activation for a series of thresholds , we can scrutinize the relative importance of various types of activated neurons.
-The percentage of various types of activated neurons for each activation threshold in the XNLI task are shown in Figure 20.
-When applying a threshold (>0), there are fewer all-shared, partial-shared, and specific neurons left in the lower layers (layers 0-10) compared to using a lower threshold (i.e., = 0).
-Under the same threshold scenario, more activated neurons appear in the second half of the model (layers 15-30).
-It is worth noting these neurons have a higher activation value than neurons activated from a lower threshold (i.e., = 0).
-Considering both the percentage of activated neurons and their corresponding activation value, it becomes apparent that the neurons in the upper layers contribute more to the output performance.
-A.5 Supplemental Results of Neuron Attribution
-A.5.1 Contribution Score of Different Tasks
-The contribution score of the four types of neurons evaluated on the XNLI task, Cross-lingual KE task, Fact Probing tasks are shown in Figures 24 - 22.
-A.5.2 Effective Score of Different Tasks
-| all-shared | partial-shared | specific | non-activated | |||||||||
-| max | min | mean | max | min | mean | max | min | mean | max | min | mean | |
-| en | 1.85 | -0.94 | 0.07 | 0.22 | -0.16 | 0.00 | 0.02 | -0.02 | 0.00 | 0.04 | -0.03 | 0.00 |
-| de | 1.03 | -0.60 | 0.02 | 0.13 | -0.13 | 0.00 | 0.07 | -0.03 | 0.00 | 0.02 | -0.01 | 0.00 |
-| es | 1.15 | -0.84 | 0.02 | 0.12 | -0.11 | 0.00 | 0.01 | -0.01 | 0.00 | 0.02 | -0.02 | 0.00 |
-| fr | 1.06 | -0.78 | 0.01 | 0.15 | -0.11 | 0.00 | 0.03 | -0.04 | 0.00 | 0.02 | -0.02 | 0.00 |
-| ru | 0.70 | -0.45 | 0.00 | 0.24 | -0.13 | 0.00 | 0.08 | -0.03 | 0.00 | 0.01 | -0.01 | 0.00 |
-| th | 0.50 | -0.90 | 0.00 | 0.17 | -0.10 | 0.00 | 0.03 | -0.05 | 0.00 | 0.01 | -0.01 | 0.00 |
-| tr | 0.82 | -0.51 | 0.03 | 0.12 | -0.12 | 0.00 | 0.04 | -0.03 | 0.00 | 0.02 | -0.02 | 0.00 |
-| vi | 0.86 | -0.68 | 0.01 | 0.15 | -0.11 | 0.00 | 0.04 | -0.04 | 0.00 | 0.02 | -0.02 | 0.00 |
-| zh | 0.52 | -0.42 | 0.00 | 0.17 | -0.20 | 0.00 | 0.08 | -0.07 | 0.00 | 0.02 | -0.01 | 0.00 |
-| pt | 1.14 | -0.83 | 0.02 | 0.11 | -0.15 | 0.00 | 0.02 | -0.02 | 0.00 | 0.02 | -0.02 | 0.00 |
-The maximum, minimum, and average effective scores of four types of neurons in 10 languages evaluated on the Cross-lingual KE (EN (edit) ALL (Test)) task are shown in Table 3.
-The maximum effective of four types of neurons across layers evaluated on the Fact Probing task, XNLI task, Cross-lingual KE (ALL (edit) EN (Test)) task are shown in Figures 25 - 27.
-A.6 Supplemental Results of Verification on Downstream Tasks
-A.6.1 Effects with LLAMA Backbone
-| accuracy | pct. | en | de | es | fr | ru | th | tr | vi | zh |
-| baseline | 0% | 59.1 | 47.6 | 50.1 | 47.0 | 49.1 | 41.4 | 40.2 | 51.6 | 46.1 |
-| w/o. all-shared | 22.42% | 3.0 | 3.6 | 4.4 | 1.9 | 4.7 | 6.9 | 3.6 | 13.5 | 4.8 |
-| w/o. partial-shared | 17.48% | 59.1 | 48.4 | 51.5 | 47.9 | 49.7 | 42.9 | 41.5 | 50.8 | 48.0 |
-| w/o. specific | 4.75% | 59.2 | 47.3 | 49.9 | 47.0 | 49.1 | 41.9 | 40.1 | 51.4 | 46.2 |
-| w/o. non-activated | 55.35% | 30.5 | 13.8 | 12.0 | 11.9 | 12.4 | 5.0 | 14.2 | 13.4 | 5.2 |
-| w/o. random | 5% | 58.7 | 47.7 | 50.2 | 48.2 | 49.0 | 41.7 | 40.0 | 49.9 | 45.7 |
-| w/o. random | 15% | 52.7 | 44.6 | 47.2 | 46.4 | 44.5 | 38.4 | 40.1 | 48.6 | 45.2 |
-| w/o. random | 25% | 46.1 | 42.4 | 41.3 | 43.3 | 40.1 | 34.5 | 39.7 | 38.7 | 40.7 |
-| w/o. random | 55% | 28.7 | 30.2 | 28.6 | 30.3 | 25.8 | 19.0 | 27.1 | 28.2 | 25.0 |
-In order to further prove the importance of all-shared neurons across LLMs, we conduct the experiments with deactivating neurons on the XNLI task with LLAMA2-7b-chat backbone.
-The results in Table 4 show that there is more significant effect when all-shared neurons are deactivated.
-It demonstrates that all-shared neurons play a key role in predicting correct answers across LLMs.